A group of farmers agree to share equally in the cost of a $48,000 piece of machinery. If they could find two more farmers to join the group, each person's share of the cost would decrease by $4000. How many farmers are presently in the group?
Answer by mananth(16946)

let x be number of farmers initially
share per farmer = 48000/x
two more join
share per farmer = 48000/(x+2)
48000/x - 48000/(x+2) = 4000
LCD = x(x+2)
multiply by LCD
48000((x+2)-48000x= 4000x(x+2)
48000x+960000-48000x=4000x^2+8000x
4000x^2+8000x-960000=0
/4000
x^2+2x-24=0
x^2+6x-4x-24=0
(x+6)(x-4)=0
x= 4 original number of farmers